# Hybrid Type-II and Type-III seesaw model for the muon $g-2$ anomaly

### v1

@@ III.1 Analytical results
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\ell}c_{L}^{\ell}\big{(}(c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2}\big{)}\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{\sqrt{2}m_{\mu}m_{F^{0}}}{8\pi^{2}m_{\delta^{-}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\nu}c_{L}^{\ell}c_{L}^{\nu}c_{R}^{\ell}\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,,$
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\ell}\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{\sqrt{2}m_{\mu}m_{F^{0}}}{8\pi^{2}m_{\delta^{-}}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\nu}\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,.$ (55)
Considering $m_{F^{\pm}}\approx m_{F^{0}}$ and $m_{\delta^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}$ we can have the following simple expressions
$a^{\textrm{NP}}_{\mu}\approx\frac{\sqrt{2}m_{\mu}m_{F}}{8\pi^{2}m_{\delta}^{2}}[x_{L}\lambda_{L}^{\ast}+\lambda_{L}x_{L}^{\ast}]s_{L}^{\nu}\cdot[-2F_{LR}^{f,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})+5F_{LR}^{S,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})]\,.$ (56)
@@ III.2 Numerical results
$a^{\textrm{NP}}_{\mu}$ $\sim\frac{m_{\mu}m_{F}}{m_{\delta}^{2}}[-\sqrt{2}\,(s^{\nu}_{L})^{2}\,\frac{m_{F}}{v_{\delta}}]\,s_{L}^{\nu}\cdot[-2\,(\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}+5\,(-\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}]$
$\sim\frac{m_{\mu}}{v_{\delta}}(s^{\nu}_{L})^{3}\sim\frac{m_{\mu}}{v_{\delta}}\left(-\sqrt{2}\,x_{L}\frac{v_{\delta}}{m_{F}}\right)^{3/2}\,,$ (57)
which can be easily seen to decrease with a certain chosen value of $x_{L}$ and increasing $m_{F}$. In the right panel of Fig. 2 we impose the $\{\lambda_{L},s^{\nu}_{L}\}$ scheme described in Eq. (II.3) as input parameters and take $\lambda_{L}=1,\,s^{\nu}_{L}=0.017$ as benchmark point. Aside from the observation that $a^{\textrm{NP}}_{\mu}$ can cover the central region of $\Delta a_{\mu}$ with $m_{F}\approx 1500\,\textrm{GeV}$ in this parameter setup, noticeable chiral enhancement is manifested by enhanced $a^{\textrm{NP}}_{\mu}$ when $m_{F}$ increases.
@@
In Fig. 4 we show $a^{\textrm{NP}}_{\mu}$ on the plane of $m_{F}$ versus $s^{\nu}_{L}$ in which the left (right) panel has $\lambda_{L}=1$ ($\lambda_{L}=0.5$). This figure can be view as the extended illustration of the right panel of Fig. 2 by making $s^{\nu}_{L}$ to vary. The dashed line is $s^{\nu}_{L}=0.017$ with respect to which the region below is allowed. We can see that in the left panel, the region with $s^{\nu}_{L}=0.017$ and $1000\,\textrm{GeV}\lesssim m_{F}\lesssim 2000\,\textrm{GeV}$ can generate $a^{\textrm{NP}}_{\mu}$ within the $1\,\sigma$ range of $\Delta a_{\mu}$. Region with smaller value of $s^{\nu}_{L}$ would require heavier $m_{F}$ and thus more significant chiral enhancement to achieve the same level of $a^{\textrm{NP}}_{\mu}$. The right panel, however, can only generate $a^{\textrm{NP}}_{\mu}$ on the edge of $1\,\sigma$ range of $\Delta a_{\mu}$ with quite heavy fermion mass $m_{F}\approx 5\,\textrm{TeV}$ with $s^{\nu}_{L}=0.017$. We also checked that all Yukawa couplings in our model satisfy the requirement of perturbativity on the shown range of Fig. 4 which can be easily seen through Eq. (II.3).
@@ III.3 Phenomenological discussions
In this section we briefly discuss the phenomenological aspects of our model. Note that the scalar triplet and VLL triplet have been scrutinized in neutrino mass generation mechanisms as Type-II [26, 27, 28, 29, 30, 31] and Type-III [32, 33] seesaw models, respectively. However, in this work we only concentrate on the physics of $(g-2)_{\mu}$ and do not require our model to produce the experimentally suggested texture of neutrino mass matrix. Instead, we require our model parameters to ensure the theoretically predicted neutrino mass to be negligibly small and close to zero. Aside from the neutrino physics, recently it has been found that the Type-II seesaw model can also address the problem of baryon asymmetry of the universe (see e.g. [59, 60, 61, 62]).
For the new physics leptons $F^{0},F^{\pm}$, Eq. (39) indicates that the VLL triplet masses are almost degenerate. Originating from the electroweak gauge interactions, $F^{0},F^{\pm}$ can be singly or pairly produced. The main decay channels of the VLL triplet are $F^{0}\rightarrow\nu h,\nu Z,\ell^{\pm}W^{\mp}$ and $F^{\pm}\rightarrow\ell^{\pm}h,\ell^{\pm}Z,\nu W^{\pm}$. If the mass spectrum satisfies $m_{F}>m_{\delta}$ which is assumed in the analysis of this work, there are additional decay channels $F^{0}\rightarrow\ell^{\pm}\delta^{\mp}$, $F^{-}\rightarrow\ell^{-}\delta^{0},\ell^{+}\delta^{--}$. Similarly, for the new physics scalars $\delta^{0},\delta^{-},\delta^{--}$, Eq. (33) indicates that the scalar triplet mass are also nearly degenerate. The scalar triplet can be pairly or singly produced through the $s$-channel mediation of SM gauge bosons $W^{\pm},\,Z$. The fermionic decay channels of the scalar triplet to SM leptons include $\delta^{0}\rightarrow\ell^{+}\ell^{-}$, $\delta^{-}\rightarrow\nu\ell^{-}$, and $\delta^{--}\rightarrow\ell^{-}\ell^{-}$. The bosonic decay channels of the scalar triplet to SM bosons include $\delta^{0}\rightarrow W^{+}W^{-}$, $\delta^{-}\rightarrow W^{-}Z$, and $\delta^{--}\rightarrow W^{-}W^{-}$, of which the partial decay widths are proportional to $v_{\delta}^{2}$. If the mass spectrum satisfies $m_{\delta}>m_{F}$, there are additional decay channels $\delta^{0}\rightarrow\ell^{\pm}F^{\mp}$, $\delta^{-}\rightarrow F^{0}\ell^{-}$, and $\delta^{--}\rightarrow F^{-}\ell^{-}$. Detailed collider analysis is beyond the scope of this paper, and recent discussions on the prospects of searches for scalar triplet around TeV at future high luminosity LHC and 100 TeV $pp$ collider can be found in [63] and [41], respectively.
@@ IV Conclusions
In this work, we investigated the muon anomalous dipole moment $a_{\mu}$ in a model that extends the SM with a scalar triplet and a vector-like lepton triplet. We identify an important ingredient overlooked in previous studies, i.e. the Yukawa interaction involving the SM Higgs doublet and the newly introduced lepton triplet. This interaction is a four-dimensional operator and thus can naturally exist. We show that this Yukawa interaction can not only induce mass mixing between leptons in the SM and the new physics sector, but also provide an additional source of chiral flipping to $a_{\mu}$. We find that there is still viable parameter space in this model to explain $\Delta a_{\mu}$, which is different from the observation reported in the existing literature. More specifically, while being consistent with the current data of neutrino mass, precision electroweak measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and fermion mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and vector-like leptons, and they can be tested at future high energy and/or high luminosity colliders.
@@ Acknowledgements.
Appendix A Reproduced results in models including only one scalar triplet or one VLL triplet
If keeping only the scalar triplet or the VLL triplet in Section III.1 by turning off relevant interactions, we can reproduce the analytical results of $a_{\mu}$ predicted in Type-II and Type-III seesaw models.
@@ Appendix A Reproduced results in models including only one scalar triplet or one VLL triplet
  • By taking $\theta^{h}=x_{L}=\lambda_{L}=0$, the $a_{\mu}^{\nu,\,\textrm{total}}$ vanishes automatically and we would reproduce the results in Type-III seesaw model as follows

### v2

@@ III.1 Analytical results
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell}c_{L}^{\ell})\big{(}\lambda_{L}^{\ast}((c_{R}^{\ell})^{2}-(s_{R}^{\ell})^{2})\big{)}]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu}c_{L}^{\ell}c^{G})(\lambda_{L}^{\ast}c_{L}^{\nu}c_{R}^{\ell}c^{G})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,,$
$\approx$ $\quad\quad\frac{m_{\mu}m_{F^{\pm}}}{4\pi^{2}m_{\delta^{--}}^{2}}\textrm{Re}[(2x_{L}s_{L}^{\ell})(\lambda_{L}^{\ast})]\cdot[-F_{LR}^{f,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})+2F_{LR}^{S,2}(\frac{m_{F^{\pm}}^{2}}{m_{\delta^{--}}^{2}})]$
$+\frac{m_{\mu}m_{F^{0}}}{4\pi^{2}m_{\delta^{-}}^{2}}\textrm{Re}[(\sqrt{2}x_{L}s_{L}^{\nu})(\lambda_{L}^{\ast})]\cdot F_{LR}^{S,2}(\frac{m_{F^{0}}^{2}}{m_{\delta^{-}}^{2}})\,.$ (65)
Utilizing $s^{\ell}_{L}\approx\sqrt{2}s^{\nu}_{L}$ from Eq. (26), as well as the mass relations $m_{\delta}\equiv m_{\delta^{0}}\approx m_{\delta^{-}}\approx m_{\delta^{--}}$ from Eq. (35) and $m_{F}\equiv m_{F^{\pm}}\approx m_{F^{0}}$ from Eq. (43), we can have the following simple expressions
$a^{\textrm{NP}}_{\mu}\approx\frac{\sqrt{2}m_{\mu}m_{F}}{4\pi^{2}m_{\delta}^{2}}\textrm{Re}[x_{L}\lambda_{L}^{\ast}]s_{L}^{\nu}\cdot[-2F_{LR}^{f,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})+5F_{LR}^{S,2}(\frac{m_{F}^{2}}{m_{\delta}^{2}})]\,.$ (66)
@@ III.2 Numerical results
$a^{\textrm{NP}}_{\mu}$ $\approx\frac{\sqrt{2}m_{\mu}m_{F}}{4\pi^{2}m_{\delta}^{2}}(\lambda_{L}x_{L})\,\sqrt{-\sqrt{2}\,x_{L}\frac{v_{\delta}}{m_{F}}}\cdot[-2\,(\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}+5\,(-\frac{1}{4})\,(\frac{m_{F}^{2}}{m_{\delta}^{2}})^{-1}]$
$\propto\frac{m_{\mu}}{m_{F}}\lambda_{L}(-x_{L})\sqrt{(-x_{L})\frac{v_{\delta}}{m_{F}}}\,,$ (67)
in which we have taken $x_{L},\lambda_{L}$ to be real numbers to simplify the expression. One can easily see that $a^{\textrm{NP}}_{\mu}$ decreases with fixed $x_{L},\lambda_{L}$ and increasing $m_{F}$. In the right panel of Fig. 2 we impose the $\{\lambda_{L},s^{\nu}_{L}\}$ scheme described in Eq. (II.3) as input parameters and take $\lambda_{L}=1,\,s^{\nu}_{L}=0.017$ as benchmark point. Aside from the observation that $a^{\textrm{NP}}_{\mu}$ can cover the central region of $\Delta a_{\mu}$ with $m_{F}\approx 1500\,\textrm{GeV}$ in this parameter setup, noticeable chiral enhancement is manifested by enhanced $a^{\textrm{NP}}_{\mu}$ when $m_{F}$ increases.
@@
In Fig. 4 we show $a^{\textrm{NP}}_{\mu}$ on the plane of $m_{F}$ versus $s^{\nu}_{L}$ in which the left (right) panel has $\lambda_{L}=1$ ($\lambda_{L}=0.5$). This figure can be view as the extended illustration of the right panel of Fig. 2 by further allowing $s^{\nu}_{L}$ to vary. The dashed line is $s^{\nu}_{L}=0.017$ with respect to which the region below is allowed. We can see that in the left panel, the region with $s^{\nu}_{L}=0.017$ and $1000\,\textrm{GeV}\lesssim m_{F}\lesssim 2000\,\textrm{GeV}$ can generate $a^{\textrm{NP}}_{\mu}$ within the $1\,\sigma$ range of $\Delta a_{\mu}$. Region with smaller value of $s^{\nu}_{L}$ would require heavier $m_{F}$ and thus more significant chiral enhancement to achieve the same level of $a^{\textrm{NP}}_{\mu}$. The right panel, however, can only generate $a^{\textrm{NP}}_{\mu}$ on the edge of $1\,\sigma$ range of $\Delta a_{\mu}$ with quite heavy fermion mass $m_{F}\approx 5\,\textrm{TeV}$ with $s^{\nu}_{L}=0.017$. We also checked that all Yukawa couplings in our model satisfy the requirement of perturbativity on the shown range of Fig. 4 which can be easily seen through Eq. (II.3).
@@ III.3 Phenomenological discussions
In this section we briefly discuss the phenomenological aspects of our model. Note that the scalar triplet and triplet have been scrutinized in neutrino mass generation mechanisms as Type-II [28, 29, 30, 31, 32, 33] and Type-III [34, 35] seesaw models, respectively. However, in this work we only concentrate on the physics of $(g-2)_{\mu}$ and do not require our model to produce the experimentally suggested texture of neutrino mass matrix. Instead, we require our model parameters to ensure the theoretically predicted neutrino mass to be negligibly small and close to zero. Aside from the neutrino physics, recently it has been found that the Type-II seesaw model can also address the problem of baryon asymmetry of the universe (see e.g. [62, 63, 64, 65]).
For the new physics leptons $F^{0},F^{\pm}$, Eq. (43) indicates that the triplet masses are almost degenerate. Originating from the electroweak gauge interactions, $F^{0},F^{\pm}$ can be singly or pairly produced. The main decay channels of the triplet are $F^{0}\rightarrow\nu h,\nu Z,\ell^{\pm}W^{\mp}$ and $F^{\pm}\rightarrow\ell^{\pm}h,\ell^{\pm}Z,\nu W^{\pm}$. If the mass spectrum satisfies $m_{F}>m_{\delta}$ which is assumed in the analysis of this work, there are additional decay channels $F^{0}\rightarrow\ell^{\pm}\delta^{\mp}$, $F^{-}\rightarrow\ell^{-}\delta^{0},\ell^{-}a^{0},\ell^{+}\delta^{--}$. Similarly, for the new physics scalars $\delta^{0},a^{0},\delta^{-},\delta^{--}$, Eq. (35) indicates that the scalar triplet mass are also nearly degenerate. The scalar triplet can be pairly or singly produced through the $s$-channel mediation of SM gauge bosons $W^{\pm},\,Z$. The fermionic decay channels of the scalar triplet to SM leptons include $\delta^{0}\rightarrow\ell^{+}\ell^{-}$,$a^{0}\rightarrow\ell^{+}\ell^{-}$, $\delta^{-}\rightarrow\nu\ell^{-}$, and $\delta^{--}\rightarrow\ell^{-}\ell^{-}$. The bosonic decay channels of the scalar triplet to SM bosons include $\delta^{0}\rightarrow W^{+}W^{-}$, $\delta^{-}\rightarrow W^{-}Z$, and $\delta^{--}\rightarrow W^{-}W^{-}$, of which the partial decay widths are proportional to $v_{\delta}^{2}$. If the mass spectrum satisfies $m_{\delta}>m_{F}$, there are additional decay channels $\delta^{0}\rightarrow\ell^{\pm}F^{\mp}$, $a^{0}\rightarrow\ell^{\pm}F^{\mp}$, $\delta^{-}\rightarrow F^{0}\ell^{-}$, and $\delta^{--}\rightarrow F^{-}\ell^{-}$. Detailed collider analysis is beyond the scope of this paper, and recent discussions on the prospects of searches for scalar triplet around TeV at future high luminosity LHC and 100 TeV $pp$ collider can be found in [43, 66, 67], respectively. The lepton triplet around TeV can also be accessible at future colliders [68, 69, 70].
@@ IV Conclusions
In this work, we investigated the muon anomalous dipole moment $a_{\mu}$ in a model that extends the SM with a scalar triplet and a lepton triplet. We identify an important ingredient overlooked in previous studies, i.e. the Yukawa interaction involving the SM Higgs doublet and the newly introduced lepton triplet. This interaction is a four-dimensional operator and thus can naturally exist. We show that this Yukawa interaction can not only induce mass mixing between leptons in the SM and the new physics sector, but also provide an additional source of chiral flip to $a_{\mu}$. We find that there is still viable parameter space in this model to explain $\Delta a_{\mu}$, which is different from the observation reported in the existing literature. More specifically, while being consistent with the current data of neutrino mass, electroweak precision measurements and the perturbativity of couplings, our model can provide new physics contribution $a_{\mu}^{\textrm{NP}}$ to cover the central region of $\Delta a_{\mu}$ with new scalar and fermion mass as low as around TeV. This mass scale is allowed by the current collider searches for doubly charged scalars and the lepton triplet, and they can be tested at future high energy and/or high luminosity colliders.
@@ Acknowledgements.
Appendix A Reproduced results in models including only one scalar triplet or one lepton triplet
If keeping only the scalar triplet or lepton triplet in Section III.1 by turning off relevant interactions, we can reproduce the analytical results of $a_{\mu}$ predicted in Type-II and Type-III seesaw models.
@@ Appendix A Reproduced results in models including only one scalar triplet or one lepton triplet
  • By taking $\theta^{h}=\theta^{a}=\theta^{G}=x_{L}=\lambda_{L}=0$, the $a_{\mu}^{\nu,\,\textrm{total}}$ vanishes automatically and we would reproduce the results in Type-III seesaw model as follows